OpenStudy (anonymous):

GEOMETRY HELP NEEDED!!! PLEASE HELP ME???? 2. The height of one square pyramid is 12 m. A similar pyramid has a height of 6 m. The volume of the larger pyramid is 400 m3. Determine each of the following, showing all your work and reasoning: a) Scale factor of the smaller pyramid to the larger pyramid in simplest form b) Ratio of the areas of the bases of the smaller pyramid to the larger pyramid c) Ratio of the volume of the smaller pyramid to the larger d) Volume of the smaller pyramid

OpenStudy (anonymous):

a) 6 m / 12 m = 1/2 b) (1/2)^2 = 1/4 c) (1/2)^3 = 1/8 d) (1/8) * 400 m^3 = 50 m^3

OpenStudy (anonymous):

Can you explain it to me? @educated

OpenStudy (anonymous):

Sure, give me a few..

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

have you covered simiilarities yet?

OpenStudy (anonymous):

what do you mean? like the lesson?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

yeah, why?

OpenStudy (jdoe0001):

well... this is a similarity issue

OpenStudy (anonymous):

i know, but the scale factor and ratio of area and volume is what is making me so confused over this problem. Do you know if educated is right?

OpenStudy (jdoe0001):

the ratio of a shape to another shape is one side length over the other to get the areas ratio, areas have 2 units, thus you'd square the same ratio to get the volume's, volumes have 3 units, so you'd cube it